# MetaphorShare: A Dynamic Collaborative Repository of Open Metaphor Datasets

### v2

@@ 1 Introduction
From the perspective of Humanities, tools for searching corpora such as Sketch Engine Kilgarriff et al. (2014), annotation tools Koller et al. (2008) and data analysis softwares have become part of the standard corpus-based research methodology. Effective adaptive support for metaphor identification and interpretation is the next step of this on-going inter-disciplinary collaboration. There were two limitations until recently: metaphor identification system were not accurate enough to be used on free text, and different researchers are looking for different metaphors.
On one side, numerous datasets labelled with metaphoric usage of words are created by researchers in many languages every year, and remain out of the scope of NLP research and evaluations of the systems because they are not publicly shared or consistently formatted. On the other, metaphor studies researchers often work on domain specific corpora, languages other than English (for which NLP has developed the most resources), and use different definitions or retain different types of metaphors for different projects, with no convenient way to systematically compare their labels with other researchers, or to rely on existing metaphor an automatic labelling tool that answer a project specific need. Moreover, these created datasets are not used by other communities or are often labeled using different schemata and formats.
We propose to help speeding up collaborations between AI/NLP communities and linguistics/metaphor studies by facilitating the unification of dataset formats and the access to existing resources by everyone, while preserving the information encoded in the original datasets. MetaphorShare is a website designed at sharing new and publicly available existing resources (see Figure 1). It is an online dynamic repository where anyone can upload or download open metaphor datasets. Our online labeling tool also offers the possibility to directly create a dataset, and optionally share it. Our hope is that the website will create a new synergy between different components of metaphor studies, limit the loss of datasets, catalog them and facilitate their automatic processing through a unified format. The targeted audience is anyone running corpus based research on metaphors, including NLP researchers for the development of metaphor processing and identification systems. [^1] An introduction to MetaphorShare is available at https://youtu.be/Fi48SOjueEE
@@ 2 Related work
Cameron and Maslen (2010) summarizes the landscape of metaphor studies. Since Lakoff and Johnson (1980) Conceptual Metaphor Theory (CMT) core role in cognitive sciences, metaphors analysis have been done differently in linguistics and literature, but also became a standard methodology in Anthropology, Educational Research, Political Science or Management research. The purpose of metaphor analysis in these disciplines is to uncover latent meaning, sometimes elusive in corpora. Recently, Baleato Rodríguez et al. (2023) introduced propaganda modeling, extending the domain of use of NLP beyond isolated metaphor processing to the perspective social media content analysis.
@@
Veale and Li (2012) develop a system of metaphor interpretation and generation, Metaphor Magnet, that relies on the harvest of stereotype from the Google n-grams. The stereotypical attributes associated to concepts are leveraged for the suggestion of relevant metaphors and their interpretation (e.g. given Google is –Microsoft, the system outputs giant with properties like lumbering and sprawling). More recently, Mao et al. (2023) release end-to-end metaphor identification and interpretation website for free text (non-domain specific): MetaPro tags metaphoric tokens or multi-word expressions in sentences, generates paraphrases, and outputs a more abstract concept mapping derived from the analysis of the sentence. The authors list limitations due to the lack of training data for inference and extended metaphors. The main difference with these two initiatives and ours is that our aim is to integrate already labeled datasets into a unified repository, encouraging the structuring of resources and expression of needs from the metaphor study community, before building automatic annotation tools. Similarly to ours, MetaPro also uses several metaphor datasets for training the models Mao et al. (2019), but in this case they do not mention more than six from the NLP community.
@@
Many open source (e.g. Label Studio, Dataturks, Doccano Nakayama et al. (2018) and Potato Pei et al. (2022)) and proprietary (e.g. LabelBox, Prodigy or Amazon Mechanical Turk) annotation tools exist. Our purpose is not to compete with these services but rather to offer a simple targeted experience for metaphor annotation, in which the output is directly mapped to the unified repository format. This way, non-expert users can avoid the need for learning a general tool, as well as having to learn how to modify given labeled outputs, which wouldn’t be trivial for most.
@@
One important usage of our website is to facilitate the creation of personalised metaphor processing models. Metaphor processing in NLP comprises many methods developed for metaphor identification Turney et al. (2011); Tsvetkov et al. (2014); Mao et al. (2019); Wachowiak and Gromann (2023), but also generation Veale (2016); Stowe et al. (2021); Chakrabarty et al. (2021b), textual Mao et al. (2018) and multimodal Kulkarni et al. (2024) interpretation, metaphor understanding through entailment Agerri et al. (2008); Chakrabarty et al. (2021a); Stowe et al. (2022), among other tasks. Ge et al. (2023) provide a comprehensive recent survey on the topic. We do not currently store images in MetaphorShare, but the extension of the database to non textual data for multimodal research is straightforward.
@@
Psycholinguistics
JANK Jankowiak (2020) 24022footnotemark: 2 50
@@
Multi-Word Expressions
@@
MIPVU
@@
Table 1: Twelve metaphor datasets included in the MetaphorShare repository. The N column shows the number of instances in the datasets. %M indicates the percentage of expressions labelled as metaphorical.
3 Data Collection and Unification
MetaphorShare aims at facilitating computational research on multiple datasets. A minimally constrained format is required to store the datasets into a database, conveniently compare the resources, search efficiently into the different fields across datasets, and run experiments on several sets at once. Metaphor datasets can be added by their authors or by anyone if their license allows it. In this section, we review the existing dataset formats and our unification choices. The website database currently contains 25 datasets. We select twelve metaphor datasets with open licenses and diverse encoded information to illustrate our unification process and as representative examples for our repository. Their characteristics are summarized in Table 1 and in more details in Appendix Table 5.
@@ 3 Data Collection and Unification
Many datasets in NLP contain labels for binary classification or paraphrasing of one delimited expression within a given context. The expressions considered might be a single token, a multi-word expression (MWE), a phrase or a compound. The task of their original release paper may be to decide if the marked expression is used metaphorically or literally, or to find a correct literal paraphrase of the metaphoric expression, among other possible analysis and tasks. The context may consist in one single word, in the case of Adjective-Noun pairs datasets, as for example in Gutiérrez et al. (2016) or Tsvetkov et al. (2014). They often consist in one sentence such as Mohammad et al. (2016) or Cardillo et al. (2010, 2017) or Jankowiak (2020). Longer contexts are sometimes directly provided, such as in Tayyar Madabushi et al. (2022b) or Haagsma et al. (2020) who include two sentences before and after the sentence containing the marked Potential Idiomatic Expression (PIE), or indirectly provided with a reference to the original document, as for example in Joseph et al. (2023) who share references to newspaper articles or Tu and Roth (2012) who link the annotated examples to sentence pointers in the British National Corpus.
@@
Steen (2010) defined an annotation guideline that is widely used in metaphor studies, the MIPVU procedure, together with the release of the VU Amsterdam Corpus (VUAC), the largest existing metaphor corpus. Other examples of English open datasets created following MIPVU, that are in MetaphorShare, are Nacey et al. (2019b) and Julich-Warpakowski (2022) on the Music Criticism domain. Many other MIPVU datasets exist in various languages with adaptations of the guideline Nacey et al. (2019a), for example Nacey (2022) for Scandinavian. In this framework, every token in a document is labelled as metaphoric or literal.
Different versions of the VUAC suiting NLP tasks, adapted for binary classification tasks Leong et al. (2018), or including additional ratings of novelty Parde and Nielsen (2018) have been released over the years. In the evaluation section of this paper, Section 5, we use a modified version of the VUAC corpus, adapted for binary classification (VUAC_BO) as described in Boisson et al. (2023).
@@
Beyond metaphorical/literal labels of expressions in context, metaphor datasets may contain more or less common additional information. There are scalar variables such as metaphoricity/figurativeness ratings Cardillo et al. (2010); Dunn (2014); Katz et al. (1988), confidence scores, concreteness scores, novelty scores Parde and Nielsen (2018), level of emotion scores Mohammad et al. (2016), frequency in corpora… They may also contain categorical variables such as PoS, textual source and genre, metaphor type, source and target concept/domains information Gordon et al. (2015). Figure 1 shows an example of the additional information encoded in the Mohammad et al. (2016) dataset for one instance, as diplayed on MetaphorShare. Some datasets, for example the LCC Mohler et al. (2016) and Gordon et al. (2015), also annotate the fragment of the context that carries the lexical information about the target of a metaphor (e.g. in the phrase ocean of happiness, the metaphoric expression is ocean and the target lexical cue of the metaphor is happiness).
@@
To simplify the search into various datasets and their handling for further system development, we require a specific input format for uploading a new resource on the website. MetaphorShare currently accepts CSV files having minimally one column named tagged_text. The column must contain text with at least one tagged expression per line. The accepted tag set is currently: < m: metaphoric, l: literal, t: target, a:semantically anomalous, u: free tag>. Each cell must minimally contain one tag and multiple tags are allowed, making possible the upload of MIPVU datasets without sentence duplication, e.g. I <m>swim</m> today in an <m>ocean</m> of <t>happiness</t>. Additional information may be added in freely named columns of the CSV file. Sentence indices can therefore be added to preserve the sentence order in a document and include metaphor in discourse datasets. Redundant fields across datasets such as PoS information have recommended names provided in the website guideline and we will see in Section 4.2 that the manual validation step by the website administrators allows a discussion with the user/author for further field names unification if needed. If any author of a dataset encounters issues to reformat a dataset into our expected format, we invite her or him to contact us by email for support.
@@ 4.1 Website architecture
Train Test set (name & size)
J&C MOH NewsMet TSV_A GUT TONG PVC MAD MAGPIE VUAC_BO
132 732 305 1645 7691 528 448 47495 3658 38323
J&C .89 .52 .68 .78 .71 .54 .61 .63 .77 .60
MOH .70 .58 .42 .73 .67 .38 .21 .59 .26 .28
NewsMet .69 .45 .61 .72 .60 .38 .33 .48 .50 .38
TSV_A .66 .50 .61 .90 .83 .52 .39 .52 .61 .51
GUT .77 .40 .66 .73 .96 .61 .76 .65 .87 .68
TONG .39 .25 .14 .0 .0 .33 .53 .50 .03 .45
PVC .56 .41 .61 .57 .52 .54 .86 .62 .78 .57
MAD .71 .45 .54 .64 .64 .52 .79 .95 .82 .58
MAGPIE .71 .45 .67 .72 .74 .56 .73 .68 .93 .64
VUAC_BO .55 .41 .65 .73 .77 .59 .55 .61 .79 .67
Table 4: Results of the cross dataset evaluation. F1-score of the metaphor class.
@@ 4.2 Uploading a new dataset
Once a dataset is formatted into a CSV file, as described in Section 3, and after its license is specified by the authors, the dataset can be uploaded by any user without registering.
@@
The required and optional fields to be filled in at uploading time are presented in Table 2. For example, if a dataset is associated with one or several publications, the references should be added in the appropriate fields, preferably in the BibTeX format. The information provided will be displayed in the dataset catalog (c.f. Figure 2 in the Appendix), and will help to further categorize the datasets in the future versions of the website.
@@
After validating the form, an immediate automatic file format check on the data is done and feedback is returned to the user for acceptance or rejection of the file. Outputs of this step can be seen in Appendix Figure 3 & 4. In a second manual validation step, the website administrators check the license information, dataset reference, and eventually suggest modifications of CSV free columns names, that contain the dataset specific additional information, in an email interaction with the user, for database field unification purposes.
@@ 4.3 Dataset Catalog
The main purpose of this page is to display the main features of each dataset included in the repository. The fields filled at uploading time, including the license chosen, are displayed in the dataset catalog page. Statistics computed from the input file are also shown and plotted, in particular the number of instances, label set and distribution. A partial view of a dataset summary in the catalog is shown in the Appendix, Figure 2.
@@ 4.5 Online data annotation
On the Edit/Label Dataset page, one can choose to directly input some text into the interface of the annotation tool, or alternatively to upload a table (in a CSV or Excel format) containing unannotated or partially annotated text. An input can then be labelled from scratch or edit labels by pairing highlighted expressions with tags (c.f. Figure 5 in the Appendix). While the MetaphorShare tagset is predefined, custom tags can also be created. The resulting labelled data (c.f. Appendix Figure 6) can then be downloaded as a CSV, or directly uploaded in MetaphorShare following the procedure described in Section 4.2. This online annotation functionality facilitates the creation of resources immediately compatible with our accepted formats, and speeds up the correction of CSV files that do not pass the automatic validation step of the dataset upload procedure. It is also an accessible tool for annotators without any computer science background.

### v3

@@ 1 Introduction
From the perspective of Humanities, tools for searching corpora such as Sketch Engine Kilgarriff et al. (2014), annotation tools Koller et al. (2008) and data analysis softwares have become part of the standard corpus-based research methodology. Effective adaptive support for metaphor identification and interpretation is the next step of this on-going inter-disciplinary collaboration. There were two limitations until recently: metaphor identification systems were not accurate enough to be used on free text, and different researchers are looking for different metaphors.
On one hand, numerous datasets labelled with metaphoric usage of words are created by researchers in many languages every year, and remain out of the scope of NLP research, tool development, and evaluations of the systems. This happens because most datasets are published in non NLP venues, they might not be publicly available, or might be tagged in an under-specified format that does not allow direct automatic processing (e.g. tagging precisely a metaphoric expression within a text). On the other hand, metaphor studies researchers often work on domain specific corpora, languages other than English - for which NLP has developed the most resources - and use different definitions or retain different types of metaphors for different projects, with no convenient way to systematically compare their labels with other researchers, or to rely on an existing metaphor automatic labelling tool that answers their project specific need.
We propose to help speeding up collaborations between AI/NLP communities and linguistics/metaphor studies by facilitating the unification of dataset formats and the access to existing resources by everyone, while preserving the information encoded in the original datasets. MetaphorShare is a website designed at sharing new and publicly available existing resources (see Figure 1). It is an online dynamic repository where anyone can upload or download open metaphor datasets. Our online labeling tool also offers the possibility to directly create a dataset, and optionally share it. Our hope is that the website will create a new synergy between different components of metaphor studies, limit the loss of datasets, catalog them, and facilitate their automatic processing through a unified format. Our target audience comprises anyone conducting corpus-based research on metaphors, including NLP researchers for the development of metaphor processing and identification systems [^1] A short introduction to MetaphorShare is available at https://youtu.be/Fi48SOjueEE.
Recent advances in metaphor processing are presented further in Section 2. The diversity of existing resources and the unified dataset input format is presented Section 3. The website is structured around four main functionalities that are described in Section 4: upload, download, search and label metaphor datasets. Section 5 shows the potential of our framework for facilitating NLP experiments, by evaluating a RoBERTa Liu et al. (2019) model on a cross-dataset classification task.
@@ 2 Related work
Cameron and Maslen (2010) summarizes the landscape of metaphor studies. The core role played by Lakoff and Johnson (1980)’s Conceptual Metaphor Theory (CMT) in Cognitive Science had an impact on the methodology developed for the analysis of metaphors in Linguistics and Literature. Metaphor analysis also became a standard approach in Anthropology, Educational Research, Political Science, or Management Research. The purpose of metaphor analysis in these disciplines is to uncover latent meaning present in discourse on a studied topic, that is sometimes conveyed elusively in corpora. Recently, Baleato Rodríguez et al. (2023) introduced propaganda modeling, integrating metaphor identification to the automatic detection of persuasive intentions.
@@
Veale and Li (2012) develop a system of metaphor interpretation and generation, Metaphor Magnet, that relies on the harvest of stereotypes from Google n-grams. The stereotypical attributes associated with concepts are leveraged for the suggestion of relevant metaphors and their interpretation (e.g. given Google is –Microsoft, the system outputs giant with properties like lumbering and sprawling). More recently, Mao et al. (2023) release an end-to-end domain independent metaphor identification and interpretation website for free text: MetaPro tags metaphoric tokens or multi-word expressions in sentences, generates paraphrases, and outputs a more abstract concept mapping derived from the analysis of the sentence. The authors list limitations due to the lack of training data for inference and extended metaphors. The main difference with these two initiatives and ours is that our aim is to integrate already labeled datasets into a unified repository, encouraging the structuring of resources and expression of needs from the metaphor study community, before building automatic annotation tools. Similarly to ours, MetaPro also uses several metaphor datasets for training the models Mao et al. (2019), but in this case they do not mention more than six datasets, all of them often used by the NLP community.
@@
Many open source (e.g. Label Studio, Dataturks, Doccano Nakayama et al. (2018) and Potato Pei et al. (2022)) and proprietary (e.g. LabelBox, Prodigy or Amazon Mechanical Turk) annotation tools exist. Our purpose is not to compete with these services but rather to offer a simple targeted experience for metaphor annotation, in which the output is directly mapped to the unified repository format. This way, non-expert users can avoid the need for learning a general tool, as well as having to learn how to modify given labeled outputs, which would not be trivial for most.
@@
One important usage of our website is to facilitate the creation of personalised metaphor processing models. Metaphor processing in NLP comprises many methods developed for metaphor identification Turney et al. (2011); Tsvetkov et al. (2014); Mao et al. (2019); Wachowiak and Gromann (2023), but also generation Veale (2016); Stowe et al. (2021); Chakrabarty et al. (2021b), textual Mao et al. (2018) and multimodal Kulkarni et al. (2024) interpretation, metaphor understanding through entailment Agerri et al. (2008); Chakrabarty et al. (2021a); Stowe et al. (2022), among other tasks. Ge et al. (2023) provide a comprehensive recent survey on the topic.
@@
Words in syntactically constrained sentences
JANK Jankowiak (2020) 240 50
@@
MWE in natural long context (NLP)
@@
Words sampled from VUAC (MIPVU)
@@
Table 1: Twelve metaphor datasets included in the MetaphorShare repository, with a single metaphor/literal tag per entry. Short contexts consist of one word or a short sentence. Long contexts correspond to multiple sentences. Datasets derived from VUAC have a document level context. The N column shows the number of instances in the datasets. %M indicates the percentage of expressions labelled as metaphorical.
[…]
Figure 2: Screenshot of the online annotation tool showing the text input area, tag selection and creation, and resulting tagged text highlighted in different colours.
3 Data Sharing
MetaphorShare aims at facilitating computational research on multiple datasets. A minimally constrained format is required to store the datasets into a database, conveniently compare the resources, search efficiently into the different fields across datasets, and run experiments on several sets at once. Metaphor datasets can be added by their authors or by anyone else if their license allows it. In this section, we review the existing dataset formats and our unification choices. The website database currently contains 25 datasets. We select twelve datasets with open licenses and diverse encoded information to illustrate our unification process and as representative examples for our repository. Their characteristics are summarized in Table 1 and in more details in the appendix (Table 4).
@@ 3 Data Sharing
Many datasets in NLP contain labels for binary classification of one delimited expression within a given context. The expressions considered might be single tokens, multi-word expressions (MWE), phrases or compounds. The task of the original data release paper may be to decide weather the marked expression is used metaphorically or literally, or to find a correct literal paraphrase of the metaphoric expression, among other possible analysis and tasks. We define a pair of constrained tags <m> and <l> to mark metaphoric of literal annotated expressions in text. Some datasets, e.g. Jankowiak (2020), also contain anomalous sentences, that are handled with an additional <a> tag.
@@
Steen (2010) defined an annotation guideline that is widely used in metaphor studies, the MIPVU procedure, together with the release of the VU Amsterdam Corpus (VUAC), the largest existing metaphor corpus. Other examples of English open datasets created following MIPVU, that are in MetaphorShare, are Nacey et al. (2019b) and Julich-Warpakowski (2022) on the Music Criticism domain. Many other MIPVU datasets exist in various languages with adaptations of the guideline Nacey et al. (2019a), for example Nacey (2022) for Norwegian. In this framework, every token in a document is labelled as metaphoric or literal.
Different versions of the VUAC suiting NLP tasks, adapted for binary classification tasks Leong et al. (2018), or including additional ratings of novelty Parde and Nielsen (2018) have been published over the years. In the evaluation section, Section 5, we use a modified version of the VUAC corpus, adapted for binary classification (VUAC_BO) as described in Boisson et al. (2023). MetaphorShare accepts entries with multiple tags per text span, as exemplified in Figure 2. MIPVU datasets are integrated in our repository with two versions fitting different possible NLP tasks: multiple tags per sentence, and one tag per sentence, with duplicated sentences.
Context.
The context of a tagged expression may consist of one single word, in the case of adjective-noun pairs datasets Gutiérrez et al. (2016); Tsvetkov et al. (2014). It often consist of one sentence such as Mohammad et al. (2016) or Cardillo et al. (2010, 2017) or Jankowiak (2020). Longer contexts are sometimes directly provided, such as in Tayyar Madabushi et al. (2022b) or Haagsma et al. (2020) who include two sentences before and after the sentence containing the marked Potential Idiomatic Expression (PIE), or indirectly provided with a reference to the original document, such as for example in Joseph et al. (2023) who share references to newspaper articles or Tu and Roth (2012) who link the annotated examples to sentence pointers in the British National Corpus.We design options to distinguish and display narrow and broad contexts surrounding a tagged expression.
@@
Beyond metaphorical/literal labels of expressions in context, metaphor datasets may contain additional information. Some datasets, for example the LCC Mohler et al. (2016) and Gordon et al. (2015), tag the fragment of the context that carries the lexical information about the target of a metaphor (e.g. in the phrase ocean of happiness, the metaphoric expression is ocean and the target lexical cue of the metaphor is happiness). We define a specific tag <t> for such lexical cues and a free tag <u> for less frequently used in-text annotations.
Additional information might also be continuous or integer variables such as metaphoricity/figurativeness ratings averages Cardillo et al. (2010); Dunn (2014); Katz et al. (1988), confidence scores, concreteness scores, novelty scores Parde and Nielsen (2018), level of emotion scores Mohammad et al. (2016), frequency in corpora… They may also contain categorical variables such as PoS, textual source and genre, metaphor type, source and target concept/domains information Gordon et al. (2015). Figure 1 shows an example of the additional information encoded in the Mohammad et al. (2016) dataset for one instance, as diplayed on MetaphorShare. Such features are unified through name recommendations in the guideline.
@@
MetaphorShare currently accepts CSV files having minimally one column named tagged_text. It must contain text with at least one tagged expression per line. The accepted tag set is: < m: metaphoric, l: literal, t: target, a:semantically anomalous, u: free tag>. Multiple tags per entry are allowed, e.g. I <m>swim</m> today in an <m>ocean</m> of <t>happiness</t>. Additional information may be added in freely named columns. Users are encouraged to follow naming recommendations for common features. Sentence indices can be added to preserve the sentence order in a document and handle datasets labeled at the discourse level. All the datasets containing examples in languages other than English could in theory be uploaded in the flexible format defined above, e.g. Polish sentences in Jankowiak (2020), Mandarin Chinese sentences in chinese-norms, Farsi sentences in levin-etal-2014-resources and Norwegian examples in Nacey (2022) can be converted smoothly. The free tag <u>, open features defined in the data file, and built-in multilingual text search functionalities of the MetaphorShare Elasticsearch database make our framework adaptable to the specificities of a language. Moreover, we will provide individual support. If any author of a dataset encounters issues reformatting a dataset into our expected format, we invite them to contact us by email for support.
@@ 4.2 Uploading a new dataset
Once a dataset is formatted into a CSV file, as described in Section 3, and once its license is specified by the authors, the dataset can be uploaded by any user.
@@
The required and optional fields to be filled in at uploading time are presented in Table 2. For example, if a dataset is associated with one or several publications, the references should be added in the appropriate fields, preferably in BibTeX format. The information provided will be displayed in the dataset catalog (c.f. Figure 4 in the appendix), and will help to further categorize the datasets in the future versions of the website.
@@
After validating the form, an immediate automatic file format check on the data is done and feedback is returned to the user for acceptance or rejection of the file. Outputs of this step can be seen in Figure 5 & 6 in the appendix. In a second manual validation step, the website administrators check the license information, dataset reference, and eventually suggest modifications of CSV free columns names, that contain the dataset specific additional information, in an email interaction with the user, for database field unification purposes.
@@ 4.3 Dataset Catalog
The purpose of this page is to display the main features of each dataset included in the repository. The fields filled at uploading time, including the license chosen, are displayed in the dataset catalog page. Statistics computed from the input file are also shown and plotted, in particular the number of instances, label set and distribution. The list of additional fields is also extracted from the input file. A partial view of a dataset summary in the catalog is shown Figure 4 in the appendix.
@@ 4.4 Search Page
Figure 3: Results of the cross dataset evaluation. F1-score of the metaphor class. Each training set contains 800 examples and the test sets sizes are shown on the $x$ axis.
@@ 4.5 Online data annotation
On the Edit/Label Dataset page, one can choose to directly input some text into the interface of the annotation tool, or alternatively to upload a table (in a CSV or Excel format) containing unannotated or partially annotated text. An input can then be labelled from scratch or edited by pairing highlighted expressions with tags (c.f. Figure 2). While the MetaphorShare tagset is predefined, custom tags can also be created. The resulting labelled data (c.f. Figure 7 in the appendix) can then be downloaded as a CSV, or directly uploaded in MetaphorShare following the procedure described in Section 4.2. This online annotation functionality facilitates the creation of resources immediately compatible with our accepted formats, and speeds up the correction of CSV files that do not pass the automatic validation step of the dataset upload procedure. It is also an accessible tool for annotators without any computer science background.